# Decoding Gray Matter: large-scale analysis of brain cell morphometry to inform microstructural modeling of diffusion MR signals

### v2

@@ 4.2 General Implications for Biophysical Modelling
  • Microglia topology is similar across different species, suggesting that a single biophysical model of diffusion in microglia would likely hold across species. In contrast, all the other cell-types likely need a dedicated model that accounts for the species-specific topological differences. Exception may be granule cells, that show small differences across species.
4.3 Lessons Learnt for Gray Matter Microstructure Imaging through dMRI
Several attempts to characterize GM microstructure through dMRI have been made in the past [73, 74, 75, 76]. These have focussed on quantifying macro and microscopic anisotropy in the GM, which are dMRI markers informative of the volume fraction and orientation distribution and dispersion of neural cell projections in the tissue. However, GM also contains cell bodies and myelinated and unmyelinated neurites that branch, curve and undulate, alongside a complex and tortuous ECS and permeable cell membranes. Modelling all these features and disentangling their concomitant contributions to the measured dMRI signal is challenging.
In recent years, some steps forward in this direction have been made thanks to the experimental observation that the typical signature of zero radius cylinders (“sticks”) representing the dMRI signal contribution from cellular projections has not been observed in GM at high b values ($>3ms/\mu m^{2}$) [77, 26]. This observation stimulated the microstructure imaging community to look for possible explanations, which led to the in-depth investigation of the most likely reasons: (i) neurite curvature [66], branching [38] and/or undulations [30]; (ii) non-negligible signal from water restricted in cell bodies [28]; (iii) non-negligible water exchange across the membrane of cellular projections [29, 30]. All of these can obscure the stick-like signature typical of the dMRI signal from water restricted in cellular projections.
[…]
The hypothesis (ii) is supported experimentally by the agreement between GM data acquired at fixed short diffusion time ($\leq$20 ms) and the soma and neurite density imaging (SANDI) model, which explicitly incorporates a soma compartment [28], and numerically by simulations in computational models of neural cells [80, 28, 39, 30]. It has been shown that SANDI estimates of soma signal fraction in the human cortex mirrors the known cytoarchitectonics of the healthy adult brain [28], and in in vivo mouse brain it strongly correlates with cell density from the Allen Brain Atlas [81]. SANDI also provides estimates of the soma MR apparent radius, but this is expected to overestimate the true radius because of volume weighting and since it is probed indirectly through the diffusivity, which is not linearly dependent on the true radius. Olesen et al. [30] estimated that this overestimation can be up to a factor 1.6 for soma radii matching those of rat pyramidal cells from [79].
The hypothesis (iii) is supported experimentally by the agreement between GM data acquired at multiple diffusion times and a modified Kärger’s model of two exchanging compartments as in both the neurite exchange imaging (NEXI) model [29] and the standard model with exchange (SMEX) [30]. The analysis of ex vivo rat brain data with SMEX suggests exchange times $t_{ex}\approx$ 5 ms, close to the $\approx$10 ms estimated by [58, 59] at 25 °C. In contrast, the analysis of in vivo rat brain data with NEXI suggests $t_{ex}\approx$20-40 ms, longer than the ex vivo case, probably in part due to the fixation process increasing membrane permeability [82]. Nevertheless, there is still no consensus regarding intra-cellular residence time in GM; estimates span values from $\approx$10 ms to considerably longer than 100 ms [57], perhaps indicating the presence of both fast and slowly exchanging components, as proposed by [83]. The estimated intra-branch and intra-domain residence times we reported (Fig.6) support this hypothesis and the necessity for including exchange between cellular projections and extracellular space for $t_{d}>20\ ms$. However, as discussed in Sec.4.1, the presence of also diffusion-mediated exchange between the cellular projections’ shaft and small protrusions on the cellular surface, such as dendritic spines, can confound and bias the estimates of permeative exchange [62, 63, 64], and caution needs to be taken when interpreting time-dependence od dMRI signals in the GM solely due to permeative exchange.
In general, exchange cannot be completely ignored, nor does exchange fully average out the restriction effects of cell membranes, supporting including both neurite exchange and a soma compartment in a parsimonious biophysical model of diffusion in GM, as proposed by Olesen et al. [30] with the SANDIX model. There are however some caveats to bear in mind when interpreting the model estimates: (a) the soma apparent MR radius is unavoidably an overestimation of the true soma radius; (b) the compartmental signal fractions reflect, apart from volume, also unknown T2 and T1 weighting, which might differ among the compartments; (c) neurite’s fast exchange could be problematic for models based on the Kärger’s model because it could violate the model’s assumptions (i.e. barrier-limited diffusion: $t_{ex}\gg d^{2}/D$, where $d$ is the neurite diameter; and $\delta\ll t_{ex}$) and lead to biased estimates; (d) molecular diffusion in structural features such as dendritic spines and glial leaflets can lead to diffusion-mediated exchange mechanisms that occur on the same time scales of permeative exchange, making the interpretation of exchange measurements through MRI solely due to permeative exchange fundamentally wrong; (e) using single diffusion encoding acquisitions it is impossible to disentangle restriction and exchange; more refined acquisitions could allow the separation and more accurate quantification of these two effects [59, 84]. The estimates of the model parameters should therefore only be taken as an indication of the true tissue features and validation against realistic numerical simulations and/or alternative measurements in controlled phantoms (e.g. biomimetic tissues and brain organoids) and/or post-mortem samples (e.g., optical, confocal or electron microscopy) remains essential.
@@ 4.4 Limitations
Our investigation, while thorough, is not exhaustive (nor could it be). Other features could certainly be measured and tabulated. However, we focus here on those deemed relevant for biophysical modeling in the current literature on microstructure imaging via dMR. We will release our code openly and freely to allow future work to complement this study with additional features and information as needed. Our focus is on a selected set of real three-dimensional reconstructions, as we aim to characterize cellular morphology in the healthy brain. Future studies can use the released code to incorporate more and improved reconstructions of healthy brain tissue as they become available (e.g., through updates to NeuroMorpho or large electron microscopy studies [85, 86]). We provide a few illustrative examples to demonstrate how this study can inform biophysical modeling of dMR signals. However, the information obtained from our investigation has broader applications, and we hope the scientific community will find it valuable for advancing our understanding of gray matter microstructure as a whole. Throughout the manuscript, we discussed the limitations of the morphometric approach used (e.g., an expected underestimation of $<20\%$ for soma volume and branch radius). Within the constraints of currently available tools, we provide reference values that were previously unavailable. We have taken great care to report error estimates and uncertainties for all measurements to account for these limitations. Future work is needed to improve morphometric algorithms, but this is beyond the scope of the current study.
@@ Acknowledgments
This work, CAR and MP are supported by the UKRI Future Leaders Fellowship MR/T020296/2.

### v4

@@ 4.2 General Implications for Biophysical Modelling
  • Microglia topology is similar across different species. Results in Fig.9 suggest that a single biophysical model of diffusion in microglia would likely hold across species. In contrast, all the other cell-types likely need a dedicated model that accounts for the species-specific topological differences.
@@ 4.3 Limitations
Whilst neural reconstructions, like the ones analysed here, offer the closest ground truth for cellular morphology they are not without their limitations. During the process of acquiring and preparing tissue samples, distortions of the cellular structure can occur, such as tissue shrinkage, or the truncation of neural projections, implicating the accuracy of the morphological measures. Furthermore, the method used to trace and reconstruct the cellular structure from microscopy images, whether done manually or automated, can impact the accuracy and detail of morphological characteristics present in the final reconstruction. As such, it is important to acknowledge that the characteristics reported here may deviate from actual in vivo measurements.
Our investigation, while thorough, is not exhaustive (nor could it be). Other features could certainly be measured and tabulated, such as spine density. However, we focus here on those deemed relevant for biophysical modeling in the current literature on microstructure imaging via dMRI. We will release our code openly and freely to allow future work to complement this study with additional features and information as needed. Our focus is on a selected set of real three-dimensional reconstructions, as we aim to characterize cellular morphology in the healthy brain. Future studies can use this code to incorporate more and improved reconstructions of healthy brain tissue as they become available (e.g., through updates to NeuroMorpho or large electron microscopy studies [81, 82]). We provide a few illustrative examples to demonstrate how this study can inform biophysical modeling of dMR signals. However, the information obtained from our investigation has broader applications, and we hope the scientific community will find it valuable for advancing our understanding of gray matter microstructure as a whole. Throughout the manuscript, we discussed the limitations of the morphometric approach used (e.g., an expected underestimation of $20\%$ for soma volume and branch radius). Within the constraints of currently available tools, we provide reference values that were previously unavailable. We have taken great care to report error estimates and uncertainties for all measurements to account for these limitations. Future work is needed to improve morphometric algorithms, but this is beyond the scope of the current study.
@@ Acknowledgments
We would like to thank Dr Kanari for informative discussion regarding the topological analysis. This work, CAR and MP are supported by the UKRI Future Leaders Fellowship MR/T020296/2. DKJ was supported by a Wellcome Trust Strategic Award (104943/Z/14/Z) and Wellcome Discovery Award (227882/Z/23/Z).
Supplementary Material of: ”Decoding Gray Matter: large-scale analysis of brain cell morphometry to inform microstructural modeling of diffusion MR signals”
Supplementary Figure 1: Table displaying the search criteria and resulting number of accepted reconstructions and corresponding rejection rate
Supplementary Figure 2: Figure demonstrating the limit depth of cellular reconstructions as a result of imaging method (some reconstructions appeared even more ‘flattened’), and the resulting FA tensors for unadjusted eigenvalues and adjusted eigenvalues. The corresponding FA for unadjusted eigenvalues = 0.67, and for adjusted eigenvalues = 0.30, so reduced dimensional representation results in increased FA.
Supplementary Figure 3: Structural feature distribution for all species
Supplementary Figure 4: Spearman’s rank correlation between structural features for all species and cell types Cell colour indicates the strength of the correlation and black cells indicate correlations that were found not to be significant (p-value adjusted according to Bonferroni correction)
Spine modeling
To estimate the impact of dendritic spines on the surface-to-volume ratio, and the resulting effects on molecular residence and exchange times, the total dendritic length for each individual cell was calculated. Based on the measured dendritic length and the prescribed spine density (1 and 2 $spine\mu m^{-1}$), the number of spines, $n$, per cell was determined. Given the surface area, $S_{spine}$, and volume, $V_{spine}$, for a single dendritic spine for each species (as reported in the literature, [83]), we added the corresponding total spine surface ($nS_{spine}$) and volume ($nV_{spine}$) to the cellular surface and volume, respectively. Finally, we recalculated the surface-to-volume ratios, including the contribution of dendritic spines, and the corresponding change in surface-to-volume ratio (Fig.5).
Supplementary Figure 5: Estimated impact of dendritic spines on surface-to-volume ratios for rodent cell types, for densities of one and two $spines/\mu m$.